# Co-Alignment: Rethinking Alignment as Bidirectional Human-AI Cognitive Adaptation

### v2

@@ 1 Introduction
The trajectory of artificial intelligence has repeatedly challenged fundamental assumptions about problem-solving and cognition. AlphaGo’s victory over Lee Sedol revealed that optimal strategies in complex domains may lie far outside human intuition, employing moves that grandmasters initially dismissed as errors but later recognized as profound innovations (silver2016mastering, ). AlphaFold’s solution to the protein folding problem—a grand challenge that resisted human efforts for decades—demonstrated that AI can navigate solution spaces in ways fundamentally different from human scientific reasoning (jumper2021highly, ). Most recently, large language models have exhibited emergent capabilities that arise not from explicit programming but from scale and self-organization, suggesting forms of intelligence that diverge from human cognitive architectures (wei2022emergent, ; bubeck2023sparks, ).
Despite these demonstrations of AI’s unique problem-solving capabilities, the dominant paradigm in AI safety and deployment remains unidirectional: we seek to align AI systems with human values, preferences, and cognitive patterns. Current alignment methods, particularly Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep, ; ouyang2022training, ), operate under three critical assumptions: (1) human preferences represent optimal or near-optimal objectives, (2) these preferences are sufficiently stable and coherent to serve as alignment targets, and (3) successful AI development means creating systems that conform to human cognitive constraints. While these assumptions may ensure short-term safety and usability, they potentially impose severe limitations on the transformative potential of artificial intelligence.
To address this challenge, we must move beyond unidirectional alignment. Consider a chess grandmaster teaming with a modern engine: peak performance arises not from the human blindly following machine lines, but from bidirectional adaptation where each partner learns from the other’s unique strengths. Current alignment methods, however, treat human cognition as a fixed constraint, leading to systems that falter out of distribution (casper2023open, ) and amplify sycophancy (sharma2024critical, ). In this work, we introduce Bidirectional Cognitive Alignment (BiCA), a framework that reconceptualizes human-AI collaboration as mutual adaptation. Drawing from cognitive science (tomasello2005understanding, ) and emergent communication (lazaridou2020emergent, ), BiCA enables agents to dynamically adjust their communication protocols and internal representations.
@@ 2 Related Work
Current AI alignment methods, dominated by Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep, ; ouyang2022training, ) and its variants like Constitutional AI (bai2022constitutional, ) and DPO (rafailov2023direct, ), assume human preferences represent optimal objectives. However, Casper et al. casper2023open identified fundamental limitations including preference instability, while Sharma et al. sharma2024critical showed that exclusive reliance on human feedback constrains AI capabilities. Traditional human-AI collaboration approaches similarly emphasize unidirectional adaptation through interactive learning (amershi2014power, ; fails2003interactive, ) and explainability (wu2022ai, ; wang2021interpretability, ). Studies reveal that effective collaboration requires mutual understanding beyond technical competence (bansal2021does, ; wang2024complementarity, ), yet existing methods maintain asymmetric relationships where only AI adapts. Recent work on scalable oversight (bowman2022measuring, ; burns2023weak, ) and cooperative IRL (hadfield2017inverse, ) begins exploring bidirectional dynamics, while safety approaches using trust regions (schulman2015trust, ; achiam2017constrained, ) inspire our KL-budget constraints for maintaining predictable behavior during adaptation.
@@
Multi-agent reinforcement learning provides foundations for collaborative interaction (zhang2021multi, ; hernandez2019survey, ), with approaches like QMIX (rashid2018qmix, ) and MADDPG (lowe2017multi, ) stabilizing training in non-stationary environments (tan1993multi, ). Mutual adaptation has been explored through co-evolution (panait2005cooperative, ), opponent modeling (he2016opponent, ), and learning with opponent-learning awareness (foerster2018learning, ), while ad hoc teamwork (stone2010ad, ; mirsky2022survey, ) addresses collaboration without prior coordination. Research on emergent communication demonstrates that agents can develop protocols through environmental pressures (lazaridou2020emergent, ; mordatch2018emergence, ), with differentiable inter-agent learning (foerster2016learning, ) enabling gradient-based optimization. Work on human-compatible protocols (lazaridou2020multi, ; andreas2017translating, ) and information-theoretic constraints (tishby2000information, ; havrylov2017emergence, ) informs our protocol generator’s use of Gumbel-Softmax sampling (jang2017categorical, ) for learning discrete yet adaptive communication based on task context.
@@
Cognitive science research on joint action (tomasello2005understanding, ; sebanz2006joint, ), theory of mind (premack1978does, ; baker2017rational, ; jara2019theory, ), and shared representations (clark1996using, ; pickering2004toward, ) provides theoretical grounding for bidirectional adaptation. Coordination without explicit communication through focal points (schelling1980strategy, ) and aligned conceptual spaces (kleiman2016coordinate, ) motivates our representation mapper, while neural synchrony findings (liu2021team, ; reinero2021inter, ) suggest biological analogs to our alignment objectives. Our instructor component draws from intelligent tutoring systems (koedinger1997intelligent, ; vanlehn2011relative, ), curriculum learning (bengio2009curriculum, ; graves2017automated, ), and zone of proximal development theory (vygotsky1978mind, ), with adaptive feedback timing (shute2008focus, ; zhu2018importance, ) informing intervention policies. Despite these advances, existing approaches suffer from fundamental limitations: unidirectional adaptation that ignores human learning potential (gabriel2020artificial, ), static rather than learned protocols (wang2023voyager, ), cognitive mismatches causing collaboration failures (srivastava2024functional, ), and poor generalization to new partners (kirk2023past, ).
@@ 4 Experiments
We validate BiCA’s effectiveness through two complementary experimental paradigms: (1) a primary collaborative navigation task (MapTalk) that tests protocol emergence and bidirectional adaptation—anchored in grounded dialogue navigation and emergent-communication setups (deVries2018TalkTheWalk, ; Thomason2020CVDN, ; Chen2019Touchdown, ; Foerster2016DIAL, )—and (2) an auxiliary latent-space exploration task (Navigator) that directly validates representation alignment via cross-model similarity and manifold/embedding alignment analyses (Kornblith2019CKA, ; Raghu2017SVCCA, ; Wang2009ManifoldAlignment, ; AlvarezMelis2018GW, ; Conneau2018MUSE, ; Harkonen2020GANSpace, ; Shen2021SeFa, ; Higgins2017BetaVAE, ). Our experimental design follows rigorous standards for reproducibility and statistical significance.

### v5

@@ 1 Introduction
The trajectory of artificial intelligence has repeatedly challenged fundamental assumptions about problem-solving and cognition. AlphaGo’s victory over Lee Sedol revealed that optimal strategies in complex domains may lie far outside human intuition, employing moves that grandmasters initially dismissed as errors but later recognized as profound innovations (silver2016mastering). AlphaFold’s solution to the protein folding problem—a grand challenge that resisted human efforts for decades—demonstrated that AI can navigate solution spaces in ways fundamentally different from human scientific reasoning (jumper2021highly). Most recently, large language models have exhibited emergent capabilities that arise not from explicit programming but from scale and self-organization, suggesting forms of intelligence that diverge from human cognitive architectures (wei2022emergent; bubeck2023sparks).
Despite these demonstrations of AI’s unique problem-solving capabilities, the dominant paradigm in AI safety and deployment remains unidirectional: we seek to align AI systems with human values, preferences, and cognitive patterns. Current alignment methods, particularly Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep; ouyang2022training), operate under three critical assumptions: (1) human preferences represent optimal or near-optimal objectives, (2) these preferences are sufficiently stable and coherent to serve as alignment targets, and (3) successful AI development means creating systems that conform to human cognitive constraints. While these assumptions may ensure short-term safety and usability, they potentially impose severe limitations on the transformative potential of artificial intelligence.
To address this challenge, we must move beyond unidirectional alignment. Consider a chess grandmaster teaming with a modern engine: peak performance arises not from the human blindly following machine lines, but from bidirectional adaptation where each partner learns from the other’s unique strengths. Current alignment methods, however, treat human cognition as a fixed constraint, leading to systems that falter out of distribution (casper2023open) and amplify sycophancy (sharma2023towards). In this work, we introduce Bidirectional Cognitive Alignment (BiCA), a framework that reconceptualizes human-AI collaboration as mutual adaptation. Drawing from cognitive science (tomasello2005understanding) and emergent communication (lazaridou2020emergent), BiCA enables agents to dynamically adjust their communication protocols and internal representations.
@@ 2 Related Work
Current AI alignment methods, dominated by Reinforcement Learning from Human Feedback (RLHF) (christiano2017deep; ouyang2022training) and its variants like Constitutional AI (bai2022constitutional) and DPO (rafailov2023direct), assume human preferences represent optimal objectives. However, Casper et al. casper2023open identified fundamental limitations including preference instability, while Sharma et al. sharma2023towards showed that exclusive reliance on human feedback constrains AI capabilities. Traditional human-AI collaboration approaches similarly emphasize unidirectional adaptation through interactive learning (amershi2014power; fails2003interactive) and explainability (wang2019interpretability; wu2022ai). Studies reveal that effective collaboration requires mutual understanding beyond technical competence (bansal2021does; vaccaro2024combinations; hemmer2025complementarity), yet existing methods maintain asymmetric relationships where only AI adapts. Recent work on scalable oversight (bowman2022measuring; burns2023weak) and cooperative IRL (hadfield2017inverse) begins exploring bidirectional dynamics, while safety approaches using trust regions (schulman2015trust; achiam2017constrained) inspire our KL-budget constraints for maintaining predictable behavior during adaptation.
@@
Multi-agent reinforcement learning provides foundations for collaborative interaction (zhang2021multi; hernandez2019survey), with approaches like QMIX (rashid2018qmix) and MADDPG (lowe2017multi) stabilizing training in non-stationary environments (tan1993multi). Mutual adaptation has been explored through co-evolution (panait2005cooperative), opponent modeling (he2016opponent), and learning with opponent-learning awareness (foerster2018learning), while ad hoc teamwork (stone2010ad; mirsky2022survey) addresses collaboration without prior coordination. Research on emergent communication demonstrates that agents can develop protocols through environmental pressures (lazaridou2020emergent; mordatch2018emergence), with differentiable inter-agent learning (foerster2016learning) enabling gradient-based optimization. Work on human-compatible protocols (lazaridou2020multi; andreas2017translating) and information-theoretic constraints (tishby2000information; havrylov2017emergence) informs our protocol generator’s use of Gumbel-Softmax sampling (jang2017categorical) for learning discrete yet adaptive communication based on task context.
@@
Cognitive science research on joint action (tomasello2005understanding; sebanz2006joint), theory of mind (premack1978does; baker2017rational; jara2019theory), and shared representations (clark1996using; pickering2004toward) provides theoretical grounding for bidirectional adaptation. Coordination without explicit communication through focal points (schelling1980strategy) and aligned conceptual spaces (kleiman2016coordinate) motivates our representation mapper, while neural synchrony findings (bevilacqua2019brain; reinero2021inter) suggest biological analogs to our alignment objectives. Our instructor component draws from intelligent tutoring systems (koedinger1997intelligent; vanlehn2011relative), curriculum learning (bengio2009curriculum; graves2017automated), and zone of proximal development theory (vygotsky1978mind), with adaptive feedback timing (shute2008focus; zhu2018importance) informing intervention policies. Despite these advances, existing approaches suffer from fundamental limitations: unidirectional adaptation that ignores human learning potential (gabriel2020artificial), static rather than learned protocols (wang2023voyager), cognitive mismatches causing collaboration failures (srivastava2024functional), and poor generalization to new partners (kirk2023past).
@@ 4 Experiments
We validate BiCA’s effectiveness through two complementary experimental paradigms: (1) a primary collaborative navigation task (MapTalk) that tests protocol emergence and bidirectional adaptation—anchored in grounded dialogue navigation and emergent-communication setups (deVries2018TalkTheWalk; Thomason2020CVDN; Chen2019Touchdown; foerster2016learning)—and (2) an auxiliary latent-space exploration task (Navigator) that directly validates representation alignment via cross-model similarity and manifold/embedding alignment analyses (Kornblith2019CKA; Raghu2017SVCCA; Wang2009ManifoldAlignment; AlvarezMelis2018GW; conneau2017word; Harkonen2020GANSpace; Shen2021SeFa; higgins2017beta). Our experimental design follows rigorous standards for reproducibility and statistical significance.
@@ 6 Conclusion
AI Agent Setup
This research was conducted through a structured human-AI collaborative framework involving multiple large language models with distinct roles. The initial limitation observation was conceived by human researchers; subsequent brainstorming and refinement were conducted in dialogue with Claude, GPT-5, and Gemini. Based on these sessions, the team generated seven candidate collaboration modes and experimental environments; after debate and automated ranking, two were jointly selected by the AI + human team. Candidate designs were then implemented and tested within the selected environments. Code prototypes were drafted primarily by GPT and Claude, with all implementations reviewed, debugged, and validated by human researchers prior to analysis. Manuscript drafting was assisted by Gemini and GPT, while final wording, methodological choices, and conclusions were determined by the authors. Orchestration followed a human-in-the-loop pattern (prompted ideation → model debate/selection → code generation → human verification), with standard research tooling (version control, experiment tracking, and reproducible scripts) and no external proprietary data.